-5/11h + 7/9 = 2/9
Solve for "h".

Answer by stanbon(75887)

(-5/11)h + 7/9 = 2/9
--------
Subtract 7/9 from both sides to get:
(-5/11)h = -5/9
-----
h = (-5/9)/(-5/11)
-----
h = 11/9
------
h = 1.222...
Cheers,
Stan H.
---------------
-----